# Beyond Baby Blues: The Child Penalty in Mental Health in Switzerland

### v1

@@ Abstract
This paper investigates the mental health penalty for women after childbirth in Switzerland. Leveraging insurance data, we employ a staggered difference-in-difference research design. The findings reveal a substantial mental health penalty for women following the birth of their first child. Approximately four years after childbirth, there is a one percentage point (p.p.) increase in antidepressant prescriptions, representing a 50% increase compared to pre-birth levels. This increase rises to 1.7 p.p. (a 70% increase) six years postpartum. The mental health penalty is likely not only a direct consequence of giving birth but also a consequence of the changed life circumstances and time constraints that accompany it, as the penalty is rising over time and is higher for women who are employed before childbirth.
@@
Keywords: Parenthood, fertility, depression, antidepressants, parental leave, staggered difference-in-difference
@@ 1 Introduction
The birth of a child is a significant life event that increases some risk factors for mental health problems. Women carry the physical burden during pregnancy and childbirth, and biological factors change. Afterwards, a child leads to more stress related to sleep deprivation, a change in the family structure and the roles and responsibilities of the parents, career uncertainties and maybe also financial pressure (Heshmati:2023). These circumstances do not only change for women but also for men. Hence, the risk of mental health problems increases. The prevalence of postpartum depression in industrialised countries for women lies between 10% and 15% (WHO:2020) and for men around 5% (Bradley:2011). This higher risk of psychiatric disorders does also not decrease if the pregnancy was planned (Munk:2015). Hence, even if a couple wants a child, it does not prevent parents from suffering from a mental health crisis.
@@
Several studies summarised by Blanchflower:2009, Clark:2008 and Dolan:2008 look at the association between parenthood and self-reported well-being and find a negative effect between having a child and happiness in the short-run. However, self-reported outcome variables might be prone to systematic under-reporting or selection bias due to lower response rates from individuals with health problems (Baetschmann:2016). Recent research by Ahammer:2023 has shed light on the different effects of having a child on mental health among men and women using drug prescription data. Women in Austria and Denmark face enduring mental health challenges after giving birth. The consumption of antidepressants significantly increases post-childbirth. Similarly, the percentage of men taking antidepressants also increases after childbirth. However, the increase is higher for women, with no significant reduction in the mental health gap between men and women over the years. This gap appears to be less related to the act of childbirth but more to the evolving roles and responsibilities that follow (Ahammer:2023).
@@
Women generally invest more time in childcare and household duties than men (Borra:2021; Guryan:2008). Consequently, they are more likely to reduce their working hours outside the home (e.g. Kleven:2020). Even when women outearn their husbands, they spend more time doing household chores (Bertrand:2015). Hence, the evolving roles and responsibilities that mothers face both at home and in the workplace can significantly impact their mental health as time pressure mediates the relationship between parenthood and mental health (Ruppanner:2019). The combination of these dual roles, along with insufficient support at home, often leads to stress and sleep deprivation, thereby increasing the risk of mental health issues for women more than for men, resulting in an increased mental health gender gap.
@@
Simple measures such as household and childcare assistance, extended time off work, or access to childcare facilities can significantly reduce the increasing mental health burden after childbirth. Therefore, family policies play a crucial role in reducing the risk of mental health issues such as depression, posttraumatic stress disorder, anxiety, and postpartum psychosis (e.g., Chatterji:2008; Avendano:2015; Aitken:2015; Hewitt:2017). A systematic review study by Heshmati:2023 suggest that an introduction of maternity leave improves mental health outcomes after childbirth as found in the majority of studies (e.g. Chatterji:2008; Avendano:2015) but the length of the leave seems to be crucial. Studies in Austria find that an increased maternity leave exacerbates the gender gap in mental health through further enforcement of traditional gender roles (e.g., Chuard:2023; Ahammer:2023). [^3] The parental leave reform extended parental leave from 24 months up to 36 months. The maximum leave for one parent was increased from 18 months to 30 months. The evidence on the effect of paternity leave on maternal mental health is scarce. Some research suggests that women’s mental health improves when partners have access to paternity leave (Bilgrami:2020) or to flexible paternity leave policies (Persson:2019).
@@
This study primarily contributes to the literature on the effect of a child on the mental health of women using anonymous prescription data and looking at medium-term outcomes (Ahammer:2023; Kravdal:2017). [^4] Ahammer:2023 use data from Austria and Denmark and Kravdal:2017 use data from Norway. We investigate the mental health penalty for women after the birth of their first child in Switzerland by using insurance data from two of the biggest Swiss insurance companies and employing a staggered difference-in-difference estimation strategy (Callaway:2021). Mental health is proxied by antidepressant prescriptions as the primary outcome. Additionally, the number of visits to psychiatrists and the number of general practitioners (GP) are used as secondary measures, as these healthcare providers issue most of these prescriptions (NZZ:2019). Switzerland has one of the highest female employment rates (BfS:2023c) in Europe, but many mothers only work part-time. This situation may be influenced by the country’s relatively short maternity leave (16 weeks), the absence of paternity leave until 2021 (extended to two weeks in 2021), and the high cost of childcare. Furthermore, Switzerland has one of the highest gender pay gaps in Europe (BfS:2023c), which may also contribute to the tendency for women to reduce their working hours more than men. As a result, many women in Switzerland face a double workload without the same level of support found in more family-friendly countries like Denmark or Norway, which may lead to different outcomes for working mothers.
@@
The findings reveal that four years after childbirth, the percentage of women having an antidepressant prescription in a specific month rises from 2% to 3%, representing a 50% increase. [^5] Women who are about to have a child, especially in the years leading up to childbirth, are less likely to have an antidepressant prescription compared to women who will never have a child or those who already have children. This percentage further climbs to 3.3% six years post-childbirth, marking a 75% increase. The results align with the findings of Ahammer:2023 for Austria and Denmark. The effect found is substantial and economically relevant as mental illnesses impose high costs on the healthcare system (Swissinfo:2008). Additionally, considering that mental health is a prerequisite for labour market participation, there is a notable loss of productivity in an economy already experiencing a shortage of skilled workers (e.g. Biasi:2021). While we do not find significant effects for visits to psychiatrists or GPs in our main specification, we observe an increase in psychiatrist visits in certain alternative specifications, suggesting some sensitivity of the results to sample choice.
@@
Last, this paper contributes to the literature on the effect of paternity leave on maternal mental health. We investigate the effect of the introduction of paternity leave in 2021 in Switzerland on the mental health of the mother. The literature on this highly relevant question is scarce. Mental health is crucial for women to enter the labour market after childbirth (e.g., Biasi:2021), and as mentioned, mental illness is costly for an economy. Therefore, economists should also consider the impact of family policies on mental health outcomes and not only on labour market outcomes (e.g., Kleven:2019; Olivetti:2017; Havnes:2011). Since the introduction was in 2021, only short-run effects can be analysed. We do not find any significant effect on mental health. Due to the low power of the analysis - only a few women have an antidepressant prescription in the first two years after childbirth; therefore, the outcome is rare - the null result should be interpreted with caution.
The remainder of the paper is structured as follows: Section 2 explains the healthcare system and provides some facts about women’s general mental health, the labour force participation and the different family policies in Switzerland. Section 3 describes the data used, how the sample has been selected for the analysis and shows some descriptives. Section LABEL:stagg_diff_diff:_empirical_strategy explains the empirical strategy, and Section LABEL:results_main presents the different results, including robustness checks, results for men and heterogeneities. The results are further discussed in Section LABEL:discussion, and potential mechanisms are explained. Last, Section LABEL:1_conclusion concludes. Additional descriptives and results are presented in Appendix LABEL:appendix:_staggered_diff_diff. Appendix LABEL:appendix:_effect_on_men presents the effect of childbirth on men’s mental health, while Appendix LABEL:appendix:_paternal_leave covers the impact of paternal leave on mothers’ mental health.
@@ 2.3 Family Policies and Labour Force
Maternity and paternity leave can help to reduce the mental health burden of mothers after childbirth. In 2005, a 16-weeks maternity leave has been introduced in Switzerland (BSV:2024). Compared to other European countries, this is a relatively short period. A two-week paternity leave has only been introduced in 2021 (BSV:2024). Furthermore, Switzerland has no flexible parental leave system that allows parents to choose who stays home with the children.
@@ 3.1 Data Sources
This paper uses data from two of the biggest Swiss insurance companies (CSS and Swica), including all incurred insurance costs for each individual over time. It is a selected sample of women insured at CSS and Swica, which corresponds to roughly 25% of the Swiss population. The dataset from CSS contains information on insured women and men born between 1971 and 2003 from the years 2010 to 2022, with information on the drugs that a person buys and the visits to GPs and psychiatrists over the years. The data from Swica contains the same information but only for women and starts in 2012. Hence, all costs related to mental health that occur for an individual are observed. Additionally, the dataset provides each insured person’s basic sociodemographic characteristics, such as age, civil status, and region of residence.
@@ 3.2 Selected Sample
The analysis focuses on women born between 1985 and 1995. We exclude older women from the sample because they are more likely to have had their first child before 2010. For example, women born in 1985 would have been 25 years old in 2010, and while many may have delayed childbearing, some are likely to have become mothers before 2010. Hence, second births would be misclassified as first births. [^9] Some births may still be misclassified, with a slightly higher potential for misclassification among individuals insured with Swica, as the data only begins in 2012 rather than 2010. Conversely, women born after 1995 are less likely to have had their first child by 2022, as those born in 1995 would only be 27 years old. Consequently, including women born after 1995 would shift the distribution of the age at the first birth to the left. Therefore, they are excluded from the analysis. [^10] These thresholds have been defined with the median age of women giving birth to their first child in Switzerland in mind, which is 31 years (BfS:2024). The sample is further refined to include only women who gave birth to their first child between 20 and 40.111111By construction, the oldest women in the sample are 37 years old when they have their first child - women born 1985 that have their first child in 2022. Hence, only the lower age is binding. This age range allows us to focus on women’s reproductive years.
@@
We further constrain the samples to the years between 2012 and 2022, a choice influenced by the introduction of the DRG on the 1st of January 2012 (DRG:2024). [^12] A Diagnostic-Related Group (DRG) is a system used by health insurance companies to categorise hospitalisation costs and determine payment amounts for hospital stays. Instead of billing for each individual service received, a fixed payment amount is predetermined based on the DRG assigned to your stay.. This significant event in the healthcare system could affect data reliability prior to 2012. After applying these criteria, the sample of women comprises $N=83,841$ individuals.
@@
To further mitigate issues regarding misclassifying first births, we use a balanced sample of women who have remained with the same insurance provider between 2012 and 2022. While this approach helps reduce the risk of bias caused by compositional changes over time, as it captures the same group of women throughout the study period, it may introduce a limitation by focusing on a specific population, as the individuals have not changed insurance for over a decade. [^13] Because women have a child at different points in time, and we only contemplate the dataset from 2012 to 2022, we still have a compositional change if using a staggered difference-in-difference approach (Callaway:2021) but a smaller one than by using an unbalanced sample. Assuming a woman has her first child in 2018, we contemplate six years before and five years after childbirth. Hence, if we look at the effect six years after childbirth, this woman is not in the sample anymore. To eliminate all compositional changes, we would have to restrict the sample to women for whom we contemplate the whole event-time of interest. This would reduce the sample size even more. Table 3.3.2 in Section 3.3.2 shows that the balanced sample only differs from the unbalanced sample in terms of nationality, language, and number of children. Thus, the balanced sample appears mostly representative of the unbalanced sample and, hence, of the broader Swiss population.
@@
This decision to use a balanced sample comes at the cost of reducing the sample size of women to $N=25,120$. Hence, the power of the analysis is reduced. An unbalanced sample is used to check for robustness. Furthermore, a balanced sample only going until 2019 ($N=18,714$) is used to avoid the potential bias of the COVID-19 pandemic on mothers’ mental health. If mothers with specific mental health choose to have a child during the pandemic and, at the same time, the pandemic has an impact on antidepressant prescriptions, this could bias the results in both directions. [^14] We do not find any change in the pattern of antidepressant prescriptions due to the Corona pandemic (see Figure LABEL:fig:descriptives_antidepri_over_years in Appendix LABEL:1_data_paternity_leave).

### v2

@@ Abstract
This paper investigates the mental health penalty for women after childbirth in Switzerland. Leveraging insurance data, we employ a staggered difference-in-difference research design. The findings reveal a substantial mental health penalty for women following the birth of their first child. Approximately four years after childbirth, there is a one percentage point (p.p.) increase in antidepressant prescriptions, representing a 50% increase compared to pre-birth levels. This increase rises to 1.7 p.p. (a 75% increase) six years postpartum. The mental health penalty is likely not only a direct consequence of giving birth but also a consequence of the changed life circumstances and time constraints that accompany it, as the penalty is rising over time and is higher for women who are employed.
@@
Keywords: antidepressants, depression, fertility, parenthood, staggered difference-in-difference,
@@ 1 Introduction
The birth of a child is a significant life event that is often associated with happiness. However, it also increases some risk factors for mental health problems. Women carry the physical burden during pregnancy and childbirth, and biological factors change. Afterwards, a child leads to more stress related to sleep deprivation, a change in the family structure and the roles and responsibilities of the parents, career uncertainties and maybe also financial pressure (Heshmati:2023). These circumstances do not only change for women but also for men. Hence, the risk of mental health problems increases. The prevalence of postpartum depression in industrialised countries for women lies between 10% and 15% (WHO:2020) and for men around 5% (Bradley:2011). This higher risk of psychiatric disorders does also not decrease if the pregnancy was planned (Munk:2015). Hence, even if a couple wants a child, it does not prevent parents from suffering from a mental health crisis. This paper examines the impact of childbirth on maternal mental health in Switzerland, using anonymized prescription data and a staggered difference-in-differences approach.
@@
Several studies summarised by Blanchflower:2009, Clark:2008 and Dolan:2008 already looked at the association between parenthood and self-reported well-being and find a negative effect between having a child and happiness in the short-run. However, self-reported outcome variables might be prone to systematic under-reporting or selection bias due to lower response rates from individuals with health problems (Baetschmann:2016). Recent research by Ahammer:2023 has shed light on the different effects of having a child on mental health among men and women using drug prescription data. Women in Austria and Denmark face enduring mental health challenges after giving birth. The consumption of antidepressants significantly increases post-childbirth. Similarly, the percentage of men taking antidepressants also increases after childbirth. However, the increase is higher for women, with no statistically significant reduction in the mental health gap between men and women over the years. This gap appears to be less related to the act of childbirth but more to the evolving roles and responsibilities that follow (Ahammer:2023). Similar effects of a decline in mental health after childbirth for women are found in Germany (Dehos:2024). They further suggest that key factors contributing to this decline may include reduced sleep, less physical activity, and limited leisure time, alongside the extensive involvement in childcare.
@@
Women generally invest more time in childcare and household duties than men (Borra:2021; Guryan:2008). Consequently, they are more likely to reduce their working hours outside the home (e.g. Kleven:2020). Even when women outearn their husbands, they spend more time doing household chores (Bertrand:2015). Hence, the evolving roles and responsibilities that mothers face both at home and in the workplace can significantly impact their mental health as time pressure mediates the relationship between parenthood and mental health (Ruppanner:2019). The combination of these dual roles, along with insufficient support at home, often leads to stress and sleep deprivation, thereby increasing the risk of mental health issues for women more than for men, resulting in an increased mental health gender gap. [^3] Another strand of the literature shows that intimate partner violance could also be associated with an increase in mental health problems after childbirth as being pregnant and having a child makes women more vulnerable and dependent on their partner (Woolhouse:2012; Chan:2022). As we are not able to add any evidence on this, we do not discuss this further.
@@
Simple measures such as household and childcare assistance, extended time off work, or access to childcare facilities can significantly reduce the increasing mental health burden after childbirth. Therefore, family policies play a crucial role in reducing the risk of mental health issues such as depression, posttraumatic stress disorder, anxiety, and postpartum psychosis (e.g., Chatterji:2008; Avendano:2015; Aitken:2015; Hewitt:2017). A systematic review study by Heshmati:2023 suggest that an introduction of maternity leave improves mental health outcomes after childbirth as found in the majority of studies (e.g. Chatterji:2008; Avendano:2015) but the length of the leave seems to be crucial. Studies in Austria find that an increased maternity leave exacerbates the gender gap in mental health through further enforcement of traditional gender roles (e.g., Chuard:2023; Ahammer:2023). [^4] The parental leave reform extended parental leave from 24 months up to 36 months. The maximum leave for one parent was increased from 18 months to 30 months. The evidence on the effect of paternity leave on maternal mental health is scarce. Some research suggests that womens’ mental health improves when partners have access to paternity leave (Bilgrami:2020) or to flexible paternity leave policies (Persson:2019).
@@
This study primarily contributes to the literature on the effect of a child on the mental health of women using anonymized prescription data and looking at medium-term outcomes (Ahammer:2023; Kravdal:2017). [^5] Ahammer:2023 use data from Austria and Denmark and Kravdal:2017 use data from Norway. We investigate the mental health penalty for women after the birth of their first child in Switzerland by using insurance data from two of the biggest Swiss insurance companies and employing a staggered difference-in-difference estimation strategy (Callaway:2021). Mental health is proxied by antidepressant prescriptions as the primary outcome. Additionally, the number of visits to psychiatrists and the number of general practitioners (GP) are used as secondary measures, as these healthcare providers issue most of these prescriptions (NZZ:2019). Switzerland has one of the highest female employment rates (BfS:2023c) in Europe, but many mothers only work part-time. This situation may be influenced by the country’s relatively short maternity leave (14 weeks), the absence of paternity leave until 2021 (extended to two weeks in 2021), and the high cost of childcare, as public kindergarten starts when a child is between three and five years depending on the region in Switzerland. Furthermore, Switzerland has one of the highest gender pay gaps in Europe (BfS:2023c), which may also contribute to the tendency for women to reduce their working hours more than men. As a result, many women in Switzerland face a double workload without the same level of support found in more family-friendly countries like Denmark or Norway, which may lead to different outcomes for working mothers.
@@
The findings reveal that four years after childbirth, the percentage of women having an antidepressant prescription in a specific month rises from 2% to 3%, representing a 50% increase. [^6] Women who are about to have a child, especially in the years leading up to childbirth, are less likely to have an antidepressant prescription compared to women who will never have a child or those who already have children. This percentage further climbs to 3.3% six years post-childbirth, marking a 75% increase. The results align with the findings of Ahammer:2023 for Austria and Denmark. The effect found is substantial and economically relevant as mental illnesses impose high costs on the healthcare system (Swissinfo:2008). Additionally, considering that mental health is a prerequisite for labour market participation, there is a notable loss of productivity in an economy already experiencing a shortage of skilled workers (e.g. Biasi:2021). While we do not find statistically significant effects for visits to psychiatrists or GPs in our main specification, we observe an increase in psychiatrist visits in certain alternative specifications, suggesting some sensitivity of the results to sample choice.
@@
Last, this paper contributes to the literature on the effect of paternity leave on maternal mental health. We investigate the effect of the introduction of paternity leave in 2021 in Switzerland on the mental health of the mother. The literature on this highly relevant question is scarce. Mental health is crucial for women to enter the labour market after childbirth (e.g., Biasi:2021), and as mentioned, mental illness is costly for an economy. Therefore, economists should also consider the impact of family policies on mental health outcomes and not only on labour market outcomes (e.g., Kleven:2019; Olivetti:2017; Havnes:2011). Since the introduction was in 2021, only short-run effects can be analysed. We do not find any statistically significant effect on mental health. Due to the low power of the analysis, only a few women have an antidepressant prescription in the first two years after childbirth; therefore, the outcome is rare, the null result should be interpreted with caution.
The remainder of the paper is structured as follows: Section 2 explains the healthcare system and provides some facts about womens’ general mental health, the labour force participation and the different family policies in Switzerland. Section 3 describes the data used, how the sample has been selected for the analysis and shows some descriptives. Section LABEL:stagg_diff_diff:_empirical_strategy explains the empirical strategy, and Section LABEL:results_main presents the different results, including robustness checks, results for men and heterogeneities. The results are further discussed in Section LABEL:discussion, and potential mechanisms are explained. Last, Section LABEL:1_conclusion concludes. Additional descriptives and results are presented in Appendix LABEL:appendix:_staggered_diff_diff. Appendix LABEL:appendix:_effect_on_men presents the effect of childbirth on mens’ mental health, while Appendix LABEL:appendix:_paternal_leave covers the impact of paternal leave on mothers’ mental health.
@@ 2.1 Healthcare System
Home visits by midwives for postnatal care and monitoring the health of both mother and newborn are covered by basic health insurance within the first 56 days postpartum. In cases of premature birth, multiple births, first-time mothers, or cesarean sections, up to 16 visits are permitted. Otherwise, the limit is 10 visits. During the first 10 days, up to five additional second daily visits are allowed, and any further visits beyond these limits require a physician’s prescription. In conclusion, the midwife visits once to twice weekly in the first two months after childbirth (BAG:2024).
@@ 2.2 Mental Health
Postpartum depression refers to depressive episodes that arise during pregnancy or within the first year after childbirth. Common symptoms include persistent sadness, loss of interest or pleasure in previously enjoyed activities, sleep disturbances, feelings of worthlessness or being a bad mother, frequent crying, and thoughts of death or suicide (APA:2020). These symptoms largely mirror those of non-postpartum depression. For mild cases of depression, psychotherapy is typically the first line of treatment. In moderate cases, psychotherapy is often combined with antidepressant medication, while in severe cases, antidepressant medication is considered essential. Depending on individual needs, additional medications such as sleep aids or anxiolytics may also be prescribed (Swissmom:2024).
@@ 2.3 Family Policies and Labour Force
Maternity and paternity leave can help to reduce the mental health burden of mothers after childbirth. In 2005, a 14-weeks maternity leave has been introduced in Switzerland. During this period, 80% of the salary is paid, and job protection extends for up to 16 weeks (BSV:2024). Compared to other European countries, this is a relatively short period. A two-week paternity leave has only been introduced in 2021 (BSV:2024). Furthermore, Switzerland has no flexible parental leave system that allows parents to choose who stays home with the children.
@@ 3.1 Data Sources
This paper uses data from two of the biggest Swiss insurance companies (CSS and Swica), including all incurred insurance costs for each individual over time. It is a selected sample of women insured at CSS and Swica, which corresponds to roughly 25% of the Swiss population. The dataset from CSS contains information on insured women and men born between 1971 and 2003 from the years 2010 to 2022, with information on the drugs that a person buys and the visits to GPs and psychiatrists over the years. The data from Swica contains the same information but only for women and starts in 2012. Hence, all costs related to mental health and childbirth that occur for an individual are observed. Additionally, the dataset provides each insured person’s basic sociodemographic characteristics, such as age, civil status, and region of residence.
@@ 3.2 Selected Sample
The analysis focuses on women born between 1985 and 1995. We exclude older women from the sample because they are more likely to have had their first child before 2010. For example, women born in 1985 would have been 25 years old in 2010, and while many may have delayed childbearing, some are likely to have become mothers before 2010. Hence, second births would be misclassified as first births. [^10] Some births may still be misclassified, with a slightly higher potential for misclassification among individuals insured with Swica, as the data only begins in 2012 rather than 2010. Conversely, women born after 1995 are less likely to have had their first child by 2022, as those born in 1995 would only be 27 years old. Consequently, including women born after 1995 would shift the distribution of the age at the first birth to the left. Therefore, they are excluded from the analysis. [^11] These thresholds have been defined with the median age of women giving birth to their first child in Switzerland in mind, which is 31 years (BfS:2024). The sample is further refined to include only women who gave birth to their first child between 20 and 40.121212By construction, the oldest women in the sample are 37 years old when they have their first child - women born 1985 that have their first child in 2022. Hence, only the lower age is binding. This age range allows us to focus on womens’ reproductive years.
@@
We further constrain the samples to the years between 2012 and 2022, a choice influenced by the introduction of the DRG on the 1st of January 2012 (DRG:2024). [^13] A Diagnostic-Related Group (DRG) is a system used by health insurance companies to categorise hospitalisation costs and determine payment amounts for hospital stays. Instead of billing for each individual service received, a fixed payment amount is predetermined based on the DRG assigned to your stay. This significant event in the healthcare system could affect data reliability prior to 2012. After applying these criteria, the sample of women comprises $N=83,841$ individuals.
@@
To further mitigate issues regarding misclassifying first births, we use a balanced sample of women who have remained with the same insurance provider between 2012 and 2022. While this approach helps reduce the risk of bias caused by compositional changes over time, as it captures the same group of women throughout the study period, it may introduce a limitation by focusing on a specific population, as the individuals have not changed insurance for over a decade. [^14] Because women have a child at different points in time, and we only contemplate the dataset from 2012 to 2022, we still have a compositional change if using a staggered difference-in-difference approach (Callaway:2021) but a smaller one than by using an unbalanced sample. Assuming a woman has her first child in 2018, we contemplate six years before and five years after childbirth. Hence, if we look at the effect six years after childbirth, this woman is not in the sample anymore. To eliminate all compositional changes, we would have to restrict the sample to women for whom we contemplate the whole event-time of interest. This would reduce the sample size even more. Specifically, individuals with poorer mental health may have less energy or motivation to change their insurance provider, leading to a sample that overrepresents those with mental health challenges. Table 3.3.2 in Section 3.3.2 shows that the balanced sample only differs from the unbalanced sample in terms of nationality, language, and number of children. Thus, the balanced sample appears mostly representative of the unbalanced sample and, hence, of the broader Swiss population.
@@
This decision to use a balanced sample comes at the cost of reducing the sample size of women to $N=25,120$. Hence, the power of the analysis is reduced. An unbalanced sample is used to check for robustness due to the reasons mentioned in the previous paragraph. Furthermore, a balanced sample only going until 2019 ($N=18,714$) is used to avoid the potential bias of the COVID-19 pandemic on mothers’ mental health. If mothers with specific mental health choose to have a child during the pandemic and, at the same time, the pandemic has an impact on antidepressant prescriptions, this could bias the results in both directions. [^15] We do not find any change in the pattern of antidepressant prescriptions due to the Corona pandemic (see Figure LABEL:fig:descriptives_antidepri_over_years in Appendix LABEL:1_data_paternity_leave).